Suppose that you are interested in the extent to which family responsibilities are related to absences from work. You measured the number of young children living at home as well as the number of work absences in the past year for a sample of 10 employees. The data is shown in the table. Using the information that you see here, calculate the value of the slope. Which of the following values comes closest to your result? Case Number of young children at home Number of days absent from work in a year 4 4 4 4 10 10 ○ +4.83 -2.09 +3.05 O +1.12
Suppose that you are interested in the extent to which family responsibilities are related to absences from work. You measured the number of young children living at home as well as the number of work absences in the past year for a sample of 10 employees. The data is shown in the table. Using the information that you see here, calculate the value of the Standard Error of the Estimate. Which of the following values comes closest to your result? Case Number of young children at Number of days absent from home work in a year 3 10 10 4.88 O5.17 O7.19 O 2.31
Answer #1
X Y X*Y X2 Y2
1 4 4 1 16
4 7 28 16 49
3 9 27 9 81
2 3 6 4 9
1 2 2 1 4
0 3 0 0 9
0 6 0 0 36
3 4 12 9 16
5 10 50 25 100
2 2 4 4 4
Sum = 21 50 133 69 324

7n. X, | = 69-212/10-249 i-1
71 Y; -324-502/10 = 74
71 Y; | 133-21 × 50/10= 28

Slope. a=-=248 SSrx241.1245

tanclard error of estimate

74 1.124528 10-2 - 2.3053